OpenStudy (anonymous):

Solve for h. V=(pi)r^2h

OpenStudy (cggurumanjunath):

V=(pi)*r^2*h

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):

V=(pi)*r^2*h divide both sides by (pi)*r^2

OpenStudy (cggurumanjunath):

r^2=r*r isn't it ? @ilyautumn

OpenStudy (cggurumanjunath):

so what is the final answer ?@ilyautumn

OpenStudy (anonymous):

h=V/pi r ^ 2

OpenStudy (cggurumanjunath):

excellent !